# Multiplicative and additive systematics in galaxy density fluctuations and clustering measurements

### v1

@@ Abstract
Galaxy clustering measurements are a key probe of the matter density field in the Universe. With the era of precision cosmology upon us, surveys rely on precise measurements of the clustering signal for meaningful cosmological analysis. However, the presence of systematic contaminants can bias the observed galaxy number density, and thereby bias the galaxy two-point statistics. As the statistical uncertainties get smaller, correcting for these systematic contaminants becomes increasingly important for unbiased cosmological analysis. We present and validate a new method for understanding and mitigating these systematics in galaxy clustering measurements (two-point function) by identifying and characterizing contaminants in the galaxy overdensity field (one-point function) using a maximum-likelihood estimator (MLE). We test this methodology with KiDS-like mock galaxy catalogs and synthetic systematic template maps. We estimate the cosmological impact of such mitigation by quantifying uncertainties and possible biases in the inferred relationship between the observed and the true galaxy clustering signal. Our method robustly corrects the clustering signal to the sub-percent level and reduces numerous additive and multiplicative systematics from $1.5\sigma$ to less than $0.1\sigma$ for the scenarios we tested. In addition, we provide an empirical approach to identifying the functional form (additive, multiplicative, or other) by which specific systematics contaminate the galaxy number density. Even though this approach is tested and geared towards systematics contaminating the galaxy number density, the methods can be extended to systematics mitigation for other two-point correlation measurements.
@@
pagerange: Multiplicative and additive systematics in galaxy density fluctuations and clustering measurementsB
@@ 1 Introduction
Systematic effects, in the context of imaging surveys, encompass a wide array of observational and instrumental effects (Sevilla-Noarbe et al., 2021) that can systematically perturb the observed galaxy overdensity field, and thereby bias the measured galaxy two-point statistics. These effects can arise from survey-dependent factors such as atmospheric and observing conditions (e.g., seeing, airmass) and detector effects, or survey-independent (astrophysical) factors such as stellar density or Galactic extinction. Accounting for all contaminating systematics can often be difficult, resulting in many possible systematic contaminants to consider (e.g., Rodríguez-Monroy et al., 2022). Moreover, for survey-independent systematics, cross-correlation of different surveys will not help detect and remove their impact, motivating a comprehensive and robust approach to their mitigation.
Multiple methods have been developed for mitigating galaxy overdensity systematics. We discuss these methods in more detail in Sec. 2.3. However, despite significant progress in mitigating galaxy overdensity systematics, challenges remain. The DES Y3 clustering results (Rodríguez-Monroy et al., 2022) are a notable illustrative example. Comparison between results from the color-selected redMaGiC sample (Rozo et al., 2016) and the flux-limited MAGLIM (Porredon et al., 2021) galaxy samples, designed for 3$\times$2-pt analysis, revealed unanticipated inconsistencies that raised questions about how well the galaxy overdensity systematics were understood and mitigated. One pressing difficulty lies in the empirical determination of the form of the systematics contamination to the overdensity field, particularly when multiple systematics effects combine in complex ways. While simulations offer valuable insights, developing empirical methods for modeling and learning about the functional form of systematics contamination is an important path forward for the field.
In this study, we introduce a new and more generalized modeling approach for systematics contamination of the galaxy overdensity field. Our approach is empirical: we provide a generalized functional form for the contamination, and a method for constraining its parameters using the real survey data. We apply this method to the mitigation of contaminants in the clustering signal and validate its performance using KiDS-like mock galaxy catalogs (Harnois-Deraps et al., 2018) with synthetic systematics incorporated into them. Our goal is to provide a method for empirically characterizing and mitigating more complex systematic contamination.
@@ 2.1 Connection to Cosmology
The matter distribution of the Universe is of fundamental interest in cosmology as it carries important information about the evolution and structure of the Universe over time. It is therefore used for testing theoretical models for cosmic acceleration and understanding the $\Lambda$CDM cosmological model. Current cosmological analyses, such as the 3$\times$2pt analysis (Krause et al., 2021; Abbott et al., 2022), use the combined information from clustering, shear, and galaxy-galaxy lensing to extract cosmological information (e.g., matter energy density, amplitude of matter fluctuations) in a Bayesian inference framework. Combining these probes is necessary because the probes are complementary and combining them enables us to marginalize over observational and theoretical uncertainties associated with single probes. We use galaxies to probe the underlying matter density field, but marginalizing over the galaxy-matter relationship requires modeling both linear and non-linear behavior, with the linear relationship that holds on large scales described as $\delta_{\text{g}}=b\delta_{m}\rightarrow w_{\text{gg}}=b^{2}w_{mm}$, where $\delta_{\text{g}/m}$ represents the galaxy/matter overdensity field, $w$ the correlation function and $b$ the linear galaxy bias.
@@ 2.3 Methods for Systematics Correction
Many methods have been developed to address the challenging problem of correcting for systematic contamination in clustering, in response to the need for robust tools to deal with contamination of potentially unknown form. For reference, Weaverdyck & Huterer (2021) provide an extensive analysis and a review of multiple decontamination methods, along with their own method using ElasticNet. Here we provide a summary of some commonly used methods, with a goal of motivating our own.
@@ 2.3.1 Mode Projection and Template Subtraction
The essence of mode projection is to remove contaminating modes at the map level. This method was introduced by Rybicki & Press (1992) and is applied to directly correct the galaxy overdensity power spectra ($C_{\ell}$) for the presence of systematic contaminants by assigning infinite variance to the systematic templates, hence modifying the signal covariance matrix. However, this method suffers from the difficulty of needing to invert the modified covariance matrix. Elsner et al. (2016) proposed an adaptation of this method that avoided this challenge. For a single contaminant template $t$ and some contamination parameter $\alpha$, the contamination model assumes that the additive terms in equation (2) can be written as $f_{\text{syst}}\!(\boldsymbol{\theta})+f^{\text{add}}_{\text{syst}}\!(\boldsymbol{\theta})/\bar{n}_{g}\!(\boldsymbol{\theta})=\alpha t(\boldsymbol{\theta})$. The mode projection formalism neglects the multiplicative term in equation (2) under the assumption that $\alpha\hat{\delta}_{g}t(\boldsymbol{\theta})\ll 1$, so the overdensity equation becomes
@@
We can write the same expression for $\sigma_{\text{tt}}^{2}$ by replacing the cross power spectrum with the auto power spectrum of the template maps. Note that in practice the covariances in equation (4) can be calculated at the pixel-level directly rather than by measuring the power spectra. The method can be straightforwardly extended to the case of multiple systematics, as well as multiple galaxy maps in the case of a tomographic analysis. Using this method, a cleaned version of the overdensity field can be obtained from equation (3), and the contamination-corrected galaxy power spectrum can be estimated from the resulting cleaned map. Note that calculating the power spectrum of the cleaned map will result in a biased estimator, and Elsner et al. (2016) details how to properly debias the estimated power spectrum of the systematics-corrected map. This method is currently implemented in the open-source galaxy clustering code NaMaster (Alonso et al., 2019).
@@ 2.3.2 Iterative Regression and ElasticNet
Finally, Weaverdyck & Huterer (2021) introduced the use of ElasticNet, a linear regression model with L1 and L2 regularization, as a mitigation method. This method assumes a contamination model from equation (2) with $f_{\text{syst}}^{\text{add}}=0$ and $f_{\text{syst}}\!(\boldsymbol{\theta})=\boldsymbol{\alpha}\bm{\mathsf{T}}\!(\boldsymbol{\theta})$, where $\boldsymbol{\alpha}$ consists of a vector of contamination parameters and $\bm{\mathsf{T}}$ is a matrix containing the systematics templates for every pixel. The contamination parameters are fit for by minimizing the loss function
@@
where $\hat{\delta}_{g}$ represents the observed galaxy overdensity, and $||\cdot||_{1}$ and $||\cdot||_{2}$ are the L1-norm and L2-norm, respectively. These two norms are meant to incentivize sparsity in the estimated parameters (L1-norm) and address possible correlations (L2-norm) between template maps. The correction to the galaxy power spectra then follows a similar procedure to that of mode projection described earlier. However, the cleaned maps are estimated by taking into account the multiplicative bias, which was neglected in previous methods. The result is a cleaned version of the overdensity map, which we call here $\hat{\delta}^{\text{clean}}_{g}$:
@@
While ElasticNet represents an advance in correcting for both additive and multiplicative systematics in the galaxy overdensity, it assumes that they are the same. This is therefore a special case of equation (2), which shows that if there are additive and multiplicative systematics in the number density, the result is different additive and multiplicative systematics in the overdensity.
@@ 2.4 Summary of Methods
The correction techniques discussed in the previous subsection have achieved notable success in reducing the impact of systematic contamination on clustering measurements and have been applied to a wide range of survey data and simulations. Most of these methods are built on several key assumptions, such as the ability to model the systematic effects accurately using a simple additive model, or one where additive and multiplicative terms are the same as each other. While methods with these assumptions have often been sufficient for existing datasets, it is by no means clear that they will be sufficient for datasets with smaller statistical uncertainties. For example, neglecting multiplicative terms (that is, $\delta_{g}$ terms in equation (2)) in the correction, or assuming they are the same as the additive term, could result in uncorrected biases in cases where those terms exist and differ from the additive systematic term.
@@
These challenges highlight the need for further advances in correction techniques and modeling that can address the more general model for systematics in galaxy overdensities shown in equation (2). This motivates us to develop a generalized and flexible approach that can accurately model and remove systematic contamination in clustering while providing some understanding of the functional form of the contamination itself. This provides a level of interpretability to the results that non-linear ML methods, such as neural networks, often lack.
@@ 3.1 Systematics Contamination Model
where $a$ and $b$ are the contamination parameters, absorbing the previous constants: $a=\alpha+{\beta}/{\bar{n}_{g}}$ and $b=\alpha$. This allows us to construct a nested model, for which the cases of purely additive ($b=0$) and purely multiplicative ($a=b$) number density contamination are special cases. Since a given systematic template might have both additive and multiplicative effects on the number density, generalization beyond these simple models is physically motivated and allows us to test our assumptions about what contamination models we use. Also note that we have chosen to consider only terms linear in $\delta_{t}$, since we do not expect contamination from higher order terms to be significant.
@@
$\hat{\delta}_{g}\!(\boldsymbol{\theta})=\delta_{g}\!(\boldsymbol{\theta})\left(1+\sum_{i}^{N_{\text{sys}}}b_{i}\delta_{t_{i}}\!(\boldsymbol{\theta})\right)+\sum_{i}^{N_{\text{sys}}}a_{i}\delta_{t_{i}}\!(\boldsymbol{\theta}).$ (12)
@@
$\hat{w}(\theta)=\langle\hat{\delta}_{g}\hat{\delta}_{g}\rangle=\langle\delta_{g}\delta_{g}\rangle\left[1+2\sum_{i}^{N_{\text{sys}}}b_{i}^{2}\langle\delta_{t_{i}}\delta_{t_{i}}\rangle\right]+\sum_{i}^{N_{\text{sys}}}a_{i}^{2}\langle\delta_{t_{i}}\delta_{t_{i}}\rangle,$ (14)
@@
$w_{\text{corr}}(\theta)=\frac{\hat{w}(\theta)-\sum_{i}^{N_{\text{sys}}}a_{i}^{2}\langle\delta_{t_{i}}\delta_{t_{i}}\rangle}{1+2\sum_{i}^{N_{\text{sys}}}b_{i}^{2}\langle\delta_{t_{i}}\delta_{t_{i}}\rangle}.$ (15)
@@ 3.2 Correlated Templates
$w_{\text{corr}}(\theta)=\frac{\hat{w}(\theta)-\sum_{i}^{N_{\text{sys}}}a_{i}^{{}^{\prime}2}\langle\delta^{{}^{\prime}}_{t_{i}}\delta^{{}^{\prime}}_{t_{i}}\rangle}{1+2\sum_{i}^{N_{\text{sys}}}b_{i}^{{}^{\prime}2}\langle\delta^{{}^{\prime}}_{t_{i}}\delta^{{}^{\prime}}_{t_{i}}\rangle},$ (16)

### v2

# Joint inference of multiplicative and additive systematics in galaxy density fluctuations and clustering measurements

@@ Abstract
Galaxy clustering measurements are a key probe of the matter density field in the Universe. With the era of precision cosmology upon us, surveys rely on precise measurements of the clustering signal for meaningful cosmological analysis. However, the presence of systematic contaminants can bias the observed galaxy number density, and thereby bias the galaxy two-point statistics. As the statistical uncertainties get smaller, correcting for these systematic contaminants becomes increasingly important for unbiased cosmological analysis. We present and validate a new method for understanding and mitigating both additive and multiplicative systematics in galaxy clustering measurements (two-point function) by joint inference of contaminants in the galaxy overdensity field (one-point function) using a maximum-likelihood estimator (MLE). We test this methodology with KiDS-like mock galaxy catalogs and synthetic systematic template maps. We estimate the cosmological impact of such mitigation by quantifying uncertainties and possible biases in the inferred relationship between the observed and the true galaxy clustering signal. Our method robustly corrects the clustering signal to the sub-percent level and reduces numerous additive and multiplicative systematics from $1.5\sigma$ to less than $0.1\sigma$ for the scenarios we tested. In addition, we provide an empirical approach to identifying the functional form (additive, multiplicative, or other) by which specific systematics contaminate the galaxy number density. Even though this approach is tested and geared towards systematics contaminating the galaxy number density, the methods can be extended to systematics mitigation for other two-point correlation measurements.
@@
pagerange: Joint inference of multiplicative and additive systematics in galaxy density fluctuations and clustering measurementsB
@@ 1 Introduction
Systematic effects, in the context of imaging surveys, encompass a wide array of observational and instrumental effects (Sevilla-Noarbe et al., 2021) that can systematically perturb the observed galaxy overdensity field, and thereby bias the measured galaxy two-point statistics. These effects can arise from survey-dependent factors such as atmospheric and observing conditions (e.g., seeing, airmass) and detector effects, or survey-independent (astrophysical) factors such as stellar density or Galactic extinction. In addition, systematics-induced modulation of galaxy redshifts across the sky can also bias our measurements of angular clustering (Baleato Lizancos & White, 2023). Accounting for all contaminating systematics can often be difficult, resulting in many possible systematic contaminants to consider (e.g., Rodríguez-Monroy et al., 2022). Moreover, for survey-independent systematics, cross-correlation of different surveys will not help detect and remove their impact, motivating a comprehensive and robust approach to their mitigation.
Multiple methods have been developed for mitigating galaxy overdensity systematics. We discuss these methods in more detail in Sec. 2.3. However, despite significant progress in mitigating galaxy overdensity systematics, challenges remain. The DES Y3 clustering results (Rodríguez-Monroy et al., 2022) are a notable illustrative example. Comparison between results from the color-selected redMaGiC sample (Rozo et al., 2016) and the flux-limited MAGLIM (Porredon et al., 2021) galaxy samples, designed for 3$\times$2-pt analysis, revealed unanticipated inconsistencies. This sparked great interest in the area of systematics treatment even though extensive validation determined the cause of the inconsistencies was most likely systematics-dependant sample selection. One pressing difficulty lies in the empirical determination of the form of the systematics contamination to the overdensity field, particularly when multiple systematics effects combine in complex ways. While simulations offer valuable insights, developing empirical methods for modeling and learning about the functional form of systematics contamination is an important path forward for the field.
In this study, we use a new and more generalized modeling approach for systematics contamination of the galaxy overdensity field. Our approach is empirical: we provide a generalized functional form for the contamination, and a method for constraining its parameters using the real survey data. We apply this method to jointly infer and mitigate contaminants in the clustering signal and validate its performance using KiDS-like mock galaxy catalogs (Harnois-Deraps et al., 2018) with synthetic systematics incorporated into them. Our goal is to provide a method for empirically characterizing and mitigating more complex systematic contamination.
@@ 2.1 Connection to Cosmology
The matter distribution of the Universe is of fundamental interest in cosmology as it carries important information about the evolution and structure of the Universe over time. It is therefore used for testing theoretical models for cosmic acceleration and understanding the $\Lambda$CDM cosmological model. Current cosmological analyses, such as the 3$\times$2pt analysis (Krause et al., 2021; Abbott et al., 2022), use the combined information from clustering, shear, and galaxy-galaxy lensing to extract cosmological information (e.g., matter energy density, amplitude of matter fluctuations) in a Bayesian inference framework. Combining these probes is advantageous because the probes are complementary and combining them enables us to marginalize over observational and theoretical uncertainties associated with single probes. We use galaxies to probe the underlying matter density field, but marginalizing over the galaxy-matter relationship requires modeling both linear and non-linear behavior, with the linear relationship that holds on large scales described as $\delta_{\text{g}}=b\delta_{m}\rightarrow w_{\text{gg}}=b^{2}w_{mm}$, where $\delta_{\text{g}/m}$ represents the galaxy/matter overdensity field, $w$ the correlation function and $b$ the linear galaxy bias.
@@ 2.3 Methods for Systematics Correction
Many methods have been developed to address the challenging problem of correcting for systematic contamination in clustering, in response to the need for robust tools to deal with contamination of potentially unknown form. For reference, Weaverdyck & Huterer (2021) provide an extensive analysis and a review of multiple decontamination methods, along with their own method using ElasticNet. For consistency we stick closely to the formulations and notation used in Weaverdyck & Huterer (2021). Here we provide a summary of some commonly used methods, with a goal of motivating our own.
@@ 2.3.1 Mode Projection and Template Subtraction
Mode projection was introduced by Rybicki & Press (1992) and is applied to directly correct the galaxy overdensity power spectra ($C_{\ell}$) for the presence of systematic contaminants by assigning infinite variance to the systematic templates, hence modifying the signal covariance matrix. However, this method suffers from the difficulty of needing to invert the modified covariance matrix. Elsner et al. (2016) proposed an adaptation of this method that avoided this challenge. For a single contaminant template $t$ and some contamination parameter $\alpha$, the contamination model assumes that the additive terms in equation (2) can be written as $f_{\text{syst}}\!(\boldsymbol{\theta})+f^{\text{add}}_{\text{syst}}\!(\boldsymbol{\theta})/\bar{n}_{g}\!(\boldsymbol{\theta})=\alpha t(\boldsymbol{\theta})$. The mode projection formalism neglects the multiplicative term in equation (2) under the assumption that $f_{\text{syst}}\delta_{g}\ll 1$, so the overdensity equation becomes
@@
We can write the same expression for $\sigma_{\text{tt}}^{2}$ by replacing the cross power spectrum with the auto power spectrum of the template maps. Note that in practice the covariances in equation (4) can be calculated at the pixel-level directly rather than by measuring the power spectra. The method can be straightforwardly extended to the case of multiple systematics, as well as multiple galaxy maps in the case of a tomographic analysis. Using this method, a cleaned version of the overdensity field can be obtained from equation (3), and the contamination-corrected galaxy power spectrum can be estimated from the resulting cleaned map. Note that calculating the power spectrum of the cleaned map will result in a biased estimator, and Elsner et al. (2016) details how to properly debias the estimated power spectrum of the systematics-corrected map. This method is currently implemented in the open-source galaxy clustering code NaMaster (Alonso et al., 2019). It is also worth noting that pseudo-Cl mode projection and ordinary least squares regression at the pixel level have been shown to be equivalent methods (Weaverdyck & Huterer, 2021).
@@ 2.3.2 Iterative Regression and ElasticNet
Finally, Weaverdyck & Huterer (2021) introduced the use of ElasticNet, a linear regression model with L1 and L2 regularization, as a mitigation method. The method uses an additive contamination model to estimate the contamination parameters for each template initially. Then, an additional map-level correction for the templates assumed to be multiplicative is done to account for multiplicative systematics. They also derive the impact of multiplicative systematics on the overdensity covariance and incorporate it into the likelihood. The likelihood for fitting the templates uses a a contamination model from equation (2) with $f_{\text{syst}}^{\text{add}}=0$ and $f_{\text{syst}}\!(\boldsymbol{\theta})=\boldsymbol{\alpha}\bm{\mathsf{T}}\!(\boldsymbol{\theta})$, where $\boldsymbol{\alpha}$ consists of a vector of contamination parameters and $\bm{\mathsf{T}}$ is a matrix containing the systematics templates for every pixel. The contamination parameters are fit for by minimizing the loss function
@@
where $\hat{\delta}_{g}$ represents the observed galaxy overdensity, and $||\cdot||_{1}$ and $||\cdot||_{2}$ are the L1-norm and L2-norm, respectively. These two norms are meant to incentivize sparsity in the estimated parameters (L1-norm) and address possible correlations (L2-norm) between template maps. The correction to the galaxy power spectra then follows a similar procedure to that of mode projection described earlier. However, as previously mentioned, the cleaned maps are estimated by taking into account the multiplicative bias for the templates assumed to have this effect, which was neglected in previous methods. The result is a cleaned version of the overdensity map, which we call here $\hat{\delta}^{\text{clean}}_{g}$:
@@
While ElasticNet represents an advance in correcting for both additive and multiplicative systematics in the galaxy overdensity, it relies on a priori information regarding which templates are multiplicative without a means for empirically choosing between the two. This becomes a special case of equation (2), which shows that if there are additive and multiplicative systematics in the number density, the result is different additive and multiplicative systematics in the overdensity. Therefore, incorrect a priori choices of which templates are multiplicative can lead to biases in the corrected overdensity map.
@@ 2.4 Summary of Methods
The correction techniques discussed in the previous subsection have achieved notable success in reducing the impact of systematic contamination on clustering measurements and have been applied to a wide range of survey data and simulations. Most of these methods are built on several key assumptions, such as the ability to treat systematics contamination as fully additive, or fully multiplicative, rather than a combination of both. While methods with these assumptions have often been sufficient for existing datasets, it is by no means clear that they will be sufficient for datasets with smaller statistical uncertainties. For example, neglecting multiplicative terms (that is, $\delta_{g}$ terms in equation (2)) in the correction, or assuming they are the same as the additive term, could result in uncorrected biases (Shafer & Huterer, 2015; Weaverdyck & Huterer, 2021) in cases where those terms exist and differ from the additive systematic term.
@@
These challenges highlight the need for further advances in correction techniques and modeling that can address the more general model for systematics in galaxy overdensities shown in equation (2). This motivates us to develop a generalized and flexible approach that can accurately model and remove systematic contamination in clustering while providing some understanding of the functional dependence on $\delta_{g}$. This provides a level of interpretability to the results that non-linear ML methods, such as neural networks, often lack.
@@ 3.1 Systematics Contamination Model
where $a$ and $b$ are the contamination parameters, absorbing the previous constants: $a=\alpha+{\beta}/{\bar{n}_{g}}$ and $b=\alpha$. This allows us to construct a nested model, for which the cases of purely additive ($b=0$) and purely multiplicative ($a=b$) number density contamination are special cases. Since a given systematic template might have both additive and multiplicative effects on the number density, generalization beyond these simple models is physically motivated and allows us to test our assumptions about what contamination models we use. Also note that we have chosen to consider only terms linear in $\delta_{t}$ since we can “feature engineer” through nonlinear transforms of one or more base templates such that the dependence is linear (Elvin-Poole et al., 2018; Weaverdyck & Huterer, 2021).
@@
$\hat{\delta}_{g}\!(\boldsymbol{\theta})\approx\delta_{g}\!(\boldsymbol{\theta})\left(1+\sum_{i}^{N_{\text{sys}}}b_{i}\delta_{t_{i}}\!(\boldsymbol{\theta})\right)+\sum_{i}^{N_{\text{sys}}}a_{i}\delta_{t_{i}}\!(\boldsymbol{\theta}).$ (12)
@@
$\hat{w}(\theta)=\langle\hat{\delta}_{g}\hat{\delta}_{g}\rangle=\langle\delta_{g}\delta_{g}\rangle\left[1+\sum_{i}^{N_{\text{sys}}}b_{i}^{2}\langle\delta_{t_{i}}\delta_{t_{i}}\rangle\right]+\sum_{i}^{N_{\text{sys}}}a_{i}^{2}\langle\delta_{t_{i}}\delta_{t_{i}}\rangle,$ (14)
@@
$w_{\text{corr}}(\theta)=\frac{\hat{w}(\theta)-\sum_{i}^{N_{\text{sys}}}a_{i}^{2}\langle\delta_{t_{i}}\delta_{t_{i}}\rangle}{1+\sum_{i}^{N_{\text{sys}}}b_{i}^{2}\langle\delta_{t_{i}}\delta_{t_{i}}\rangle}.$ (15)
@@ 3.2 Correlated Templates
$w_{\text{corr}}(\theta)=\frac{\hat{w}(\theta)-\sum_{i}^{N_{\text{sys}}}a_{i}^{{}^{\prime}2}\langle\delta^{{}^{\prime}}_{t_{i}}\delta^{{}^{\prime}}_{t_{i}}\rangle}{1+\sum_{i}^{N_{\text{sys}}}b_{i}^{{}^{\prime}2}\langle\delta^{{}^{\prime}}_{t_{i}}\delta^{{}^{\prime}}_{t_{i}}\rangle},$ (16)